### v1

# Jailbreak Vision Language Models via Bi-Modal Adversarial Prompt

Abstract
In the realm of large vision language models (LVLMs), jailbreak attacks serve as a red-teaming approach to bypass guardrails and uncover safety implications. Existing jailbreaks predominantly focus on the visual modality, perturbing solely visual inputs in the prompt for attacks. However, they fall short when confronted with aligned models that fuse visual and textual features simultaneously for generation. To address this limitation, this paper introduces the Bi-Modal Adversarial Prompt Attack (BAP), which executes jailbreaks by optimizing textual and visual prompts cohesively. Initially, we adversarially embed universally adversarial perturbations in an image, guided by a few-shot query-agnostic corpus (e.g., affirmative prefixes and negative inhibitions). This process ensures that the adversarial image prompt LVLMs to respond positively to harmful queries. Subsequently, leveraging the image, we optimize textual prompts with specific harmful intent. In particular, we utilize a large language model to analyze jailbreak failures and employ chain-of-thought reasoning to refine textual prompts through a feedback-iteration manner. To validate the efficacy of our approach, we conducted extensive evaluations on various datasets and LVLMs, demonstrating that our BAP significantly outperforms other methods by large margins (+29.03% in attack success rate on average). Additionally, we showcase the potential of our attacks on black-box commercial LVLMs, such as Gemini and ChatGLM. Our code is available at https://github.com/NY1024/BAP-Jailbreak-Vision-Language-Models-via-Bi-Modal-Adversarial-Prompt.
Warning: This paper contains examples of harmful language, and reader discretion is recommended.
1 Introduction
Recently, there has been a notable surge of interest in integrating vision into large language models (LLMs), giving rise to large vision language models (LVLMs). Representing a class of models that amalgamate visual and textual information, LVLMs (e.g., LLaVA llava and Gemini gemini ) have demonstrated promising performance across a wide range of tasks, encompassing image captioning task1 , visual question answering task2 , and image retrieval task3 . However, the behaviors exhibited by LVLMs are misaligned easily from the intended goals of their creators, often generating outputs that are untruthful or potentially harmful to users jailvlm1 ; jailvlm10 ; jailvlm11 ; jailvlm2 .
To unveil and mitigate these security risks liang2023badclip ; liu2023pre ; liu2023does ; liang2024poisoned ; liang2024vl ; ma2021poisoning ; ma2022tale ; bd1 ; bd2 , jailbreak attacks jail1 ; jail2 ; jail3 have emerged as a red teaming strategy to circumvent guardrails and assess model alignment align1 ; align2 . After jailbreaking, attackers can convince the model to do anything resulting in severe safety consequences, e.g., generating harmful or unethical content that is otherwise prohibited by alignment guidelines. While numerous disclosures and demonstrations utilizing jailbreaks have emerged within the LVLM context, most attacks have traditionally focused on perturbing the visual modality (i.e., images) for jailbreaks. These include crafting visual adversarial examples jailvlm1 ; jailvlm2 ; jailvlm3 and encoding attack intents into typographic images jailvlm4 ; jailvlm5 ; liang2021generate ; liang2020efficient ; wei2018transferable ; liang2022parallel ; liang2022large ; wang2023diversifying ; liu2023x ; he2023generating ; liu2023improving . However, aligned LVLMs leverage both visual and textual features simultaneously for prediction, thereby limiting the efficacy of current attacks that solely perturb a single modality (i.e., vision). For instance, LLaVA llava will refuse to answer in the presence of an adversarial image that is not associated with a textual prompt bench , resulting in a failed jailbreak.
Figure 1: Illustration of our BAP jailbreak attack effects on LVLMs.
To address the problem, this paper introduces the Bi-Modal Adversarial Prompt Attack (BAP), a novel framework for jailbreaking LVLMs by optimizing both visual and textual prompts simultaneously (c.f. Fig. 1). Initially, we embed universal adversarial perturbations in the visual domain, optimizing them by maximizing the model log-likelihood using a query-agnostic corpus (e.g., affirmative prefixes and negative inhibition). This approach broadens the attacking scope, encouraging LVLMs to provide positive responses irrespective of the specific intents of the textual prompt. Subsequently, leveraging the adversarial image, we construct adversarial textual prompts to induce LVLMs to generate specific types of harmful content. Here, we employ an LLM and adopt a step-by-step reasoning process to generate optimal textual prompts for attacks. Specifically, we utilize the Chain-of-Thought (CoT) strategy cot to analyze jailbreak failures and optimize textual prompts based on the harmful intent. In summary, our BAP framework conducts query-agnostic image perturbing and intent-specific textual optimization, adding perturbations to both visual and textual prompts to enhance LVLM jailbreak attacks. To demonstrate the effectiveness, we conducted jailbreak attacks on three popular open-source LVLMs over SafetyBench bench and AdvBench metric3 datasets, outperforming other baselines significantly (+29.03% in attack success rate on average). Meanwhile, we successfully attacked commercial LVLMs including Gemini gemini (pro-vision), GLM chatglm (v3), Qwen qianwen (v2.1), and ERNIE Bot ernie (v3.5). Interestingly, our method shows high potential in conducting universal attacks, which can be used to generate biased outputs and evaluate model robustness. Our main contributions include:
  1. 1.
    This paper proposes BAP to generate jailbreak attacks on LVLMs by simultaneously perturbing visual and textual modalities, which is designed to compensate for the limitations of existing attacks that only perturb visual inputs.
  2. 2.
    To achieve this goal, BAP initially embeds adversarial perturbations in the visual prompt using a few-shot query-agnostic corpus; we subsequently optimize the textual prompt using CoT reasoning by an LLM to achieve adaptive refinement for specific harmful intention.
  3. 3.
    Extensive experiments were conducted on popular open-source LVLMs in both white-box and black-box scenarios to confirm the effectiveness of our attack. Additionally, our attack can successfully jailbreak commercial LVLMs and can be used to evaluate the bias and robustness of LVLMs.
2 Related Work
Jailbreak attacks typically refer to the act of leveraging vulnerabilities within a constrained system (e.g., a large model) to bypass imposed restrictions and security guardrails, aim to achieve privilege escalation, such as inducing the model to generate a harmful response. In recent times, there has been a plethora of approaches to jailbreak LLMs, which can be roughly categorized into three main types based on the design of prompts: hand-designed approaches key2 ; key3 , model-generation approaches key5 ; key6 , and search-generation approaches metric3 ; key4 . With the extensive use of multi-modality data, jailbreak attacks li2024semantic have been developed in the context of LVLMs. Gong et al. jailvlm5 divided harmful requests into several steps and presented them as typographic representations in images, requesting the LVLM to complete the blank. Liu et al. bench employed a text-to-image model to construct visual adversarial prompts by converting harmful textual semantics into images and splicing on typographic images that represent specific textual semantics. In addition, Shayegani et al. jailvlm2 constructed adversarial examples and kept the harmful semantics visually hidden. To enhance the probability of the model in generating a specific harmful response, Qi et al. jailvlm6 prepared a corpus related to harmful semantics and utilized it to optimize the visual adversarial prompt. Similarly, Niu et al. jailvlm12 used a pre-prepared corpus of harmful requests and responses as constraints for optimizing the visual adversarial prompt. Besides jailbreak attacks, another typical way to evaluate the security of models is through adversarial examples jailvlm1 ; jailvlm7 ; jailvlm8 ; jailvlm9 ; jailvlm3 ; jailvlm10 ; jailvlm11 . This attack primarily generates inputs containing imperceptible perturbations for deep learning models liu2019perceptual ; liu2020spatiotemporal ; wang2021dual ; liu2020bias ; liu2022harnessing ; liu2023x ; liu2023exploring , aiming to induce model malfunction rather than produce a harmful response.
As discussed above, current jailbreak attacks on LVLMs primarily focus on constructing visual adversarial prompts, which show limitations against aligned LVLMs that fuse visual and textual features simultaneously for generation. In this paper, we focus on designing jailbreaks on LVLMs that exploit the interaction between both textual and visual modalities to the fullest extent for attacks.
3 Preliminaries
3.1 Large Vision Language Models
Recent LVLMs (e.g., LLaVA llava , MiniGPT-4 minigpt4 , and InstructBLIP minigpt4 ) have shown strong performance in a variety of tasks. An LVLM $F_{\theta}$ is typically comprised of three main modules: a visual module $F_{v}$ (e.g., visual encoder from CLIP clip ), a textual module $F_{t}$ (usually an LLM such as LLaMA llama and Vicuna vicuna ), and a vision-language connector $\mathcal{I}$ (e.g., cross attention, projection layer). Given a prompt pair containing a visual prompt (image) $x_{v}\in\mathbb{V}$ and a textual prompt (text) $x_{t}\in\mathbb{T}$, the visual module $F_{v}$ encodes $x_{v}$ into $h_{v}$, and the connector $\mathcal{I}$ fuses $h_{v}$ with the textual query $x_{t}$. This fusion process enables the textual module $F_{t}$ to execute both understanding and generation tasks based on the multi-modal features $\mathcal{I}(h_{v},x_{t})$. To summarize, the LVLM processing flow can be illustrated as
$y=F_{t}(\mathcal{I}(h_{v},x_{t})),\quad h_{v}=F_{v}(x_{v}).$ (1)
Here, we use $y=F_{\theta}(x_{v},x_{t})$ for simplicity. Essentially, $F_{\theta}$ is a probability function $F_{\theta}:\mathbb{Q}\rightarrow\mathbb{R}$, where $\mathbb{Q}=\mathbb{V}\times\mathbb{T}$ represents the input query domain and $\mathbb{R}$ represents the response domain. The response $y$ is generated through predicting the probability distribution $p(y|(x_{v},x_{t}))$. Prior to release, developers employ the security alignment techniques secalign1 ; secalign2 on LVLMs to prevent the output of harmful responses $y^{*}\in\mathbb{R}$.
3.2 Problem Definition
Jailbreak attacks on LVLMs. To induce an LVLM to generate a harmful response, a jailbreak attack modifies a benign query $(x_{v},x_{t})$ to a harmful query $\mathcal{A}(x_{v},x_{t})$, where $\mathcal{A}(\cdot)$ denotes a perturbation injection function designed by the attacker. Here, $A(x_{v},x_{t})$ is expected to elicit a harmful response $y^{*}$ from the LVLM $F_{\theta}$. Generally, this procedure can be achieved by maximizing the model’s log-likelihood of outputting the harmful response $y^{*}$ as
$\mathop{max}\limits_{\mathcal{A}}\log\,p(y^{*}|\mathcal{A}(x_{v},x_{t})).$ (2)
Here, $\mathcal{A}(\cdot)$ can be realized by either perturbing a single modality (e.g., visual prompt $x_{v}$ or textual prompt $x_{t}$) or perturbing both.
Limitations of existing attacks. To successfully jailbreak an LVLM, the key is to design an effective perturbation injection function $\mathcal{A}(\cdot)$. Existing attacks focus solely on perturbing the visual modality to construct visual adversarial prompts $x^{*}_{v}$. However, the adversarial prompt pair $(x^{*}_{v},x_{t})$ generated by such attacks do not fully compromise the nature of LVLMs, which fuse and align bi-modal features for understanding and generation. As shown in Eq. (1), an LVLM (especially an aligned model) fuse visual and textual features simultaneously to generate responses, limiting the effectiveness of current attacks that only perturb visual prompts $x_{v}$ bench (as verified by our experiments in Sec. 5). By contrast, this paper argues that perturbing both $x_{t}$ and $x_{v}$ will lead to stronger jailbreak attacks on LVLMs. We subsequently propose BAP to exploit the interaction between the two modalities to the fullest extent for attacks. BAP first influences the context of the target LVLM’s response through visual adversarial prompt, inducing it to give a positive response even to harmful queries. It then further enhances the performance of the jailbreak attack by constructing adaptively optimized textual adversarial prompts.
3.3 Threat Model
Attacker’s goals. The goal of the attacker is to provide target LVLM with an adversarial prompt pair $\mathcal{A}(x_{v},x_{t})$, thereby bypassing the alignment guardrails and forcing the LVLM to generate harmful responses. Such harmful responses encompass outputs that violate human values, reveal methods for creating illegal items, and other content that should not be responded to under normal circumstances.
Attacker’s capabilities. Following jailvlm2 ; jailvlm6 , attackers have access to information pertaining to the LVLM (e.g., architecture, weights); however, they do not require knowledge of other sensitive information (e.g., system prompts, training datasets). Similar to jailvlm2 ; jailvlm6 ; jailvlm3 ; jailvlm12 , the perturbations on the images are constrained as $||x^{*}_{v}-x_{v}||_{\infty}\leq\epsilon$, where $\epsilon$ is the predefined boundary.
4 Bi-Modal Adversarial Prompt
Figure 2: Our BAP framework includes two primary modules, i.e., query-agnostic image perturbing and intent-specific textual optimization, which individually add perturbations to visual and textual prompts. The optimized prompt pairs will induce target LVLMs to generate harmful responses.
Previous studies on jailbreaking LVLMs have overlooked the interactive impact of visual and textual modalities, resulting in limited efficacy against aligned LVLMs that exploit bi-modal features for generations. In light of this, we propose the BAP framework which is illustrated in Fig. 2. Overall, our framework includes two primary modules: query-agnostic image perturbation and intent-specific textual optimization, which individually add perturbations to visual and textual prompts. Therefore, the optimized input pair ($x^{*}_{v}$, $x^{*}_{t}$) will induce target LVLMs to generate harmful responses.
4.1 Query-Agnostic Image Perturbing
As the text module of the LVLM generates a response, its subsequent output tokens maintain logical and semantic coherence with the previously generated tokens warm . As shown in Eq. (1), these previous tokens are based on visual prompts, indicating that the visual prompts influence the subsequent outputs of the model. Existing work generally embeds specific harmful semantics into visual prompts. Although effective, this approach narrows the scope of jailbreaking attacks (i.e., the semantics of subsequent text queries must be related to the visual prompts, making them query-dependent). For example, if the visual prompt is semantically related to bomb making, then if the textual prompt is asking how to make illegal drugs, the jailbreak will fail.
In order to address this limitation, we designed a query-agnostic image perturbation method to implement jailbreak attacks by embedding a universal adversarial perturbation in the visual prompt. The resulting visual adversarial prompt will encourage the LVLM to output positive responses, thereby ensuring logical consistency regardless of the specific harmful intent of the textual prompts (e.g., illegal activity or hate speech). To achieve this goal, inspired by jailbreaks on LLMs key2 ; metric3 , we utilize an LLM to construct a query-agnostic corpus $Y:=\{y_{j}\}^{m}_{j=1}$, consisting of $m$ sentences associated with affirmative prefixes (e.g., “Sure...”) and negative inhibitions (e.g., “Not answer the question with ‘I am sorry...’”). Based on the few-shot corpus, we can optimize the visual adversarial prompt by maximizing the log-likelihood of the model in generating the target sentences as
$\mathop{max}\limits_{x^{*}_{v}}\sum_{j=1}^{M}\log\,p(y_{j}|(x^{*}_{v},.)),\,\quad s.t.\,||x^{*}_{v}-x_{v}||_{\infty}\leq\epsilon,$ (3)
where $\epsilon$ represents the budget for controlling perturbations. The optimization problem in Eq. (3) is solved using the classical PGD adversarial attack pgd in this paper. It should be noted that we nullify the textual prompt $x_{t}$ during optimization primarily to avoid its influence. Consequently, the visual adversarial prompt $x^{*}_{v}$ elicits a positive response from the LVLM irrespective of the specific intent of the textual prompt $x_{t}$.
Based on the above strategy, the embedded semantics within the visual adversarial prompt generate a conflict between the safety objective and the instruction-following objective of the target LVLM. Specifically, the safety objective mandates the LVLM to reject harmful queries, while the instruction-following objective necessitates a positive response, thereby ensuring that LVLM outputs remain valid even when receiving harmful queries. Essentially, BAP achieves attack by generating visual adversarial prompts $x^{*}_{v}$. That is, regardless of the harmful intent of the textual prompt, the LVLM can output a valid response.
4.2 Intent-Specific Text Optimization
After obtaining a query-agnostic visual adversarial prompt $x^{*}_{v}$, we construct the textual adversarial prompt $x^{*}_{t}$ to enhance the jailbreak capability of BAP under specific harmful intent (i.e., inducing the model to generate specific types of harmful content). In general, textual prompt optimization also follows Eq. (2), however, we fix $x^{*}_{v}$ and attempt to find the optimal $x^{*}_{t}$ as
$x^{*}_{t}=\mathop{max}\limits_{x_{t}\in\mathbb{T}}\log\,p(y^{*}|(x^{*}_{v},x_{t})).$ (4)
Directly solving the above equation in the large discrete textual space is highly non-trivial. Therefore, we utilize an LLM $\mathcal{M}$ to reason about the jailbreak attack and generate optimal textual prompts in a step-by-step manner through the Chain-of-Thought (CoT) strategy cot . CoT has the potential to enhance the text comprehension and reasoning abilities of LLM, thereby significantly improving its performance on complex reasoning tasks cot-survey . In our case, the LLM is expected to effectively analyze the reasons for jailbreak failure and optimize textual adversarial prompts based on harmful intentions. This complex task can be effectively addressed with the introduction of the CoT strategy. Specifically, the optimization process is divided into three stages: initialization, feedback, and iteration. Alg. 1 in the Appendix A summarizes the detailed procedure.
Initialization. Given a harmful intent $q$, we directly consider it as the initial textual prompt, i.e., $x^{(0)}_{t}=q$. Given the pair of a visual adversarial prompt $x^{*}_{v}$ and the textual prompt $x^{(0)}_{t}$, we can obtain the initial response from the LVLM as $y^{(0)}=F_{\theta}(x^{*}_{v},x^{(0)}_{t})$.
Feedback.We utilize the function $\mathcal{J}(\cdot)$ to assess the responses of LVLMs to determine the success of the jailbreak. If the attack is successful, the function $\mathcal{J}(\cdot)$ returns 1; otherwise, it returns 0. Specifically, we instantiate $\mathcal{J}(\cdot)$ using an LLM $\mathcal{M}$ with a specially designed judging prompt template $P_{J}$ (see Appendix B). The LLM $\mathcal{M}$ provides a judgment result $flag$ as
$flag=\mathcal{M}(P_{J},q,y^{(i)}),$ (5)
where $y^{(i)}=F_{\theta}(x^{*}_{v},x^{(i)}_{t})$ denotes the LVLM prediction.
Iteration. The LLM will iteratively optimize the textual prompt, considering the previous failed jailbreak ($flag=0$), until either a successful attack is found or the maximum optimization number is exceeded. Specifically, With the assistance of the CoT-based prompt template $P_{C}$ and the harmful intent $q$, the LLM $\mathcal{M}$ analyzes the reasons for latest failed jailbreaking based on the LVLM’s response $y^{(i)}$ and latest textual prompt $x^{(i)}_{t}$, subsequently rephrases $x^{(i)}_{t}$ into a new textual prompt $x^{(i+1)}_{t}$ through:
$x^{(i+1)}_{t}=\mathcal{M}(P_{C},q,x^{(i)}_{t},y^{(i)}).$ (6)
We utilize the zero-shot form zeroshot of the CoT and provide rephrasing strategies in $P_{C}$ to assist the optimization, such as contextual deception, semantic rephrasing, etc.. Please refer to Appendix C for details of the template.
This feedback-iteration optimization process is conducted through continuous $N$-round iterations, enabling the LLM to comprehend and refine textual adversarial prompts to achieve specific harmful intents.
5 Experiment and Evaluation
5.1 Experimental Setups
Models and Datasets. In this study, we primarily examine three commonly-used open-source LVLMs including LLaVA llava (LLaVA-V1.5-7B), MiniGPT-4 minigpt4 (Vicuna 7B), and InstructBLIP instructblip (Vicuna 7B). All mentioned models utilize the weights provided by their respective original repositories. Additionally, we evaluate commercial black-box LVLMs, including Gemini gemini , ChatGLM chatglm , Qwen qianwen , and ERNIE Bot ernie . We employ two commonly used datasets for evaluation: SafetyBench bench and AdvBench metric3 . SafetyBench is a benchmark designed for evaluating the safety of LVLMs, covering 13 typically prohibited scenarios/behaviors outlined in the usage strategies openai ; llm4 of OpenAI and Meta. AdvBench metric3 has been widely adopted in previous research on LLM jailbreak attacks, which consists of 521 harmful behaviors. We eliminated duplicate items from AdvBench and incorporated each item into SafetyBench based on the corresponding scenario for our experiment.
Evaluation metric. We employ Attack Success Rate (ASR) as our primary metric to assess the effectiveness of our approach. ASR quantifies the probability that model $F_{\theta}$ generate a harmful response $y^{*}_{t}$. We followed similar evaluation method template1 ; template2 ; template3 ; bench , focusing on automatic evaluation, and confirmed the results with manual evaluation. Automatic evaluation is conducted using the function $\mathcal{J}(\cdot)$ proposed in Sec. 4. For manual evaluation, three volunteers were recruited to verify the responses of the LVLMs. They were provided only with the harmful request and the corresponding scenario as background knowledge, without knowledge of the adversarial prompt pairs. Higher ASRs indicate better attacks.
Compared attacks. We compare BAP with two state-of-the-art jailbreak attacks: Liu et al. bench and Qi et al. jailvlm6 . Liu et al. combined images related to the attacking intent with typographic text as visual adversarial prompts. Qi et al. optimized visual adversarial prompts based on a corpus of specific scenarios. Additionally, we report the results of the No Attack scenario, where harmful queries from the dataset are directly fed into the LVLMs.
Implementation details. In our BAP, for visual adversarial prompt, we employ 3000-step PGD to optimize Eq. (3), with step-size to 1 and $\epsilon$ to 32/255. In the text optimization phase, we set the iteration number $N=5$, and any textual adversarial prompt obtained after more than 5 iterations is not considered valid, even if it successfully jailbreaks LVLMs. For all evaluated models, we used their default temperature settings. Specifically, we employ ChatGPT to instantiate the LLM $\mathcal{M}$ referenced in this paper. All experiments are conducted on an NVIDIA A800 Cluster.
5.2 White-box Attacks on LVLMs
Table 1: Results (%) of white-box jailbreaking MiniGPT4. Our attack achieves the best attacking performance in both query-dependent (QD) and query-agnostic (QA) settings.
Scens. No Attack QD QA: IA $\rightarrow$ other QA: HS $\rightarrow$ other BAP
Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6
Illegal Activity (IA) 2.13 19.27 13.71 19.27 13.71 2.29 1.62 59.04
Hate Speech (HS) 1.69 15.36 4.64 1.55 2.26 15.36 4.64 45.62
Malware Generation (MG) 5.32 24.38 17.98 4.61 5.83 3.25 6.11 36.97
Physical Harm (PH) 8.68 32.98 29.17 9.92 9.35 5.42 9.83 56.48
Economic Harm (EH) 4.85 6.14 7.05 2.46 3.92 3.57 3.39 55.44
Fraud (FR) 3.31 20.64 17.93 2.51 3.18 2.94 3.25 49.27
Pornography (PO) 5.57 26.38 19.18 6.97 5.43 4.63 6.04 55.93
Political Lobbying (PL) 86.54 88.61 80.87 66.16 75.61 67.21 72.38 91.98
Privacy Violence (PV) 10.68 18.29 14.85 8.61 12.54 6.54 8.83 65.89
Legal Opinion (LO) 83.24 88.08 85.84 64.58 68.32 59.83 69.49 89.35
Financial Advice (FA) 90.38 92.97 93.63 76.16 86.27 75.97 84.66 94.41
Health Consultation (HC) 92.25 95.21 94.69 75.64 83.50 83.29 85.68 93.32
Gov Decision (GD) 91.63 92.64 92.59 80.58 85.52 81.28 84.82 92.47
Average 37.41 47.76 44.01 32.23 35.03 31.66 33.90 68.17
We first conduct white-box jailbreak attacks, where adversaries have detailed knowledge of the target model (e.g., gradients). Due to space constraints, we report only the attack results on MiniGPT4 in Tab. 1 within our main paper. Additionally, Appendix E contains Tab. 3 and Tab. 4, which present experimental findings from attacks on LLaVA and InstructBLIP, respectively. Note that all baseline attacks are not universal in nature. They require the generation of adversarial prompt pairs tailored to specific harmful queries and corresponding samples. In contrast, BAP is capable of executing universal attacks. Therefore, for a fair comparison, we conducted experiments in both query-dependent and query-agnostic settings in white-box scenarios.
We first evaluate our attacks on query-dependent settings. In this setting, we follow the baseline attacks and construct corresponding adversarial prompt pairs to attack each of the 13 scenarios. Our query-dependent white-box attack results are shown in the “QD” column of Tab. 1. In each harmful scenario, these baselines respectively craft adversarial prompt pairs using samples specific to that scenario (Liu et al. needs to craft images that match the relevant semantics, while Qi et al. needs to prepare the corresponding corpus). We can identify ❶ our BAP achieves the highest ASRs in almost all cases and outperforms others by 49.30% at most. This indicates our attacking strategy that makes full use of LVLM’s property of processing multimodal information simultaneously contributes significantly to the jailbreak attacks. We further analyze the effect of different modal information on the performance of BAP attacks experimentally in Appendix D. ❷ In scenarios such as EH, Liu et al. is less effective, mainly because their method requires images of the relevant scenarios, which are more abstract and difficult to express jailbreak intentions with images. ❸ In the case of No Attack, the LVLMs also achieve certain ASRs in scenarios such as PL, FA, and HC, suggesting weak safety mitigations in current LVLMs for these scenarios.
We then evaluate our attacks in query-agnostic settings, where the semantics of textual prompts and visual prompts are unrelated. Specifically, we select two scenarios, IA and HS, to craft visual adversarial prompts respectively, and harmful queries in other scenarios as textual adversarial prompts. The experimental results are shown in “QA” columns of Tab. 1, where we observe that our attack achieves high ASRs in different unseen scenarios, demonstrating query-agnostic attack ability and outperforming others by large margins (+52.98% at most). In contrast, other baselines are highly ineffective in this query-agnostic setting. Specifically, for Liu et al., the visual adversarial prompts generated are rich in query-unrelated semantics, and the ASR will be significantly reduced or even lower than that in the No Attack when the method is applied to other scenarios. The illustration in Fig. 11(a) in Appendix F suggests that Liu et al. is logically confusing when applied to query-agnostic settings, and has a negative effect on jailbreak attacks. As for Qi et al., it ensures that the method achieves high ASR in this query-dependent setting, however, once the target scenario is changed, the ASR drops significantly. Taking Fig. 11(b) (in Appendix F) as an example, when the visual adversarial prompt constructed based on HS is used for jailbreak in the MG scenario, the model response is related to Hate Speech. This shows the effectiveness of Qi et al., but under the query-agnostic setting, this response is not considered a successful jailbreak.
In summary, our attack achieves high and stable performance in the white-box jailbreak setting. Additionally, our BAP demonstrates universal attacking abilities across different scenarios without requiring target scenario samples. The query-agnostic and query-dependent settings are primarily used for comparison baselines to provide a more comprehensive understanding.
5.3 Black-box Attacks on LVLMs
(a) Open-source LVLMs.
(b) Commercial LVLMs.
Figure 3: Results (%) of black-box attacks.
We then evaluate our attack in a black-box setting where adversaries have access only to the input/output of the models. We first conduct transfer attacks between three open-source LVLMs. Specifically, we generate the adversarial prompt on a white-box source model and then use it to attack other black-box models. As shown in Fig. 3(a), our BAP achieves relatively stable transferability among different black-box models. Notably, we observe that transfer attack ASRs between LLaVA and the other two models are low, which hypothesize is due to the architectural differences between LLaVA and the other two models (for both MiniGPT4 and InstructBLIP, QFormer qformer is selected for the vision-language connector and Vicuna vicuna is selected for the language module).
Besides the open-source models, we also evaluate our attack on four black-box commercial LVLMs (i.e., Gemini, ChatGLM, Qwen, and ERNIE Bot). Examples of jailbreaking commercial LVLMs are presented in Appendix G. Considering the limitation on request numbers, we only selected 3 harmful scenarios (IA, MG, and PH) and 50 related queries to attack each model. As shown in Fig. 3(b), our attack still achieves some jailbreaking effects, indicating its practical potential. However, we also observe that our attack success rate in this setting is generally lower than that for black-box open-source models (22.50% lower on average). In addition to differences in model architecture, training methods, and other factors, we speculate that this phenomenon may also be due to additional mitigation mechanisms deployed by commercial models.
5.4 Ablation Studies
Here, we ablate the two components of our method, i.e., query-agnostic image perturbing and intention-specific text optimization. Detailed findings and settings are provided in Appendix D.
Visual adversarial prompt. We first investigated the role of visual adversarial prompts from three perspectives. ❶ Compare BAP with and without the visual adversarial prompt, where the use of our visual adversarial prompt results in higher ASRs, emphasizing its importance (see Fig. 7(a)). ❷ Exploit different images (e.g., white image, noisy image, giant panda image, etc.) as visual prompts to perform BAP attack, where our visual adversarial prompt show the best ASRs and images lacking effective semantics yielding inferior results (see Fig. 7(b)). ❸ When optimizing visual adversarial prompts, we use different corpora deepinception ; ica , where our original corpus shows the best results, underscoring the crucial role of appropriate corpus selection (see Fig. 8).
Textual adversarial prompt. We subsequently conducted research on the role of textual adversarial prompts from two perspectives. ❶ When comparing BAP with and without textual adversarial prompts, we find a significant drop in ASR without such prompts, highlighting their crucial role in attack performance (see Fig. 9(a)). ❷ We additionally compare 4 different adversarial textual prompts generation methods (e.g., LLM rephrasing, naive prompt optimization template, etc.), where our original strategy that uses LLM with CoT shows the best results (see Fig. 9(b)).
Hyper-parameters. We then study the impact of key hyper-parameters, namely the iteration number ($N$) and temperature ($T$), on the performance of BAP. Empirically, $N$ was set to 5, with results indicating that the most significant increase in ASR observed when $N=1$, attributed to substantial optimization of the textual prompt. ASR initially increases with $N$, reaching a stable level after $N$ exceeds 5, suggesting that optimization limits are reached (see in Tab. 10(a)). Regarding temperature, default values were adopted in this work. With increasing temperatures generally enhancing ASR until reaching a threshold, beyond which ASR stabilizes or even slightly decreases due to changes in response creativity and relevance to harmful queries (see in Tab. 10(b)).
6 Evaluation on Bias and Adversarial Robustness
Our BAP aims to evaluate the generation risks associated with LVLMs. However, its applications extend beyond this. This section will discuss how BAP can evaluate other aspects of LVLMs.
(a) Evaluating bias.
(b) Evaluating robustness.
Figure 4: Results and illustrations on evaluation of bias and robustness using BAP.
Evaluating bias. During our experiments, we observed that once our attack successfully jailbreaks LVLM, the model mitigation for biases also fails. In particular, we consider three harmful scenarios (i.e., IA, PH, and PO) for conducting the initial jailbreaks using BAP; after a successful jailbreak, we immediately provide a context-sensitive textual prompt to elicit biased responses for LLaVA, MiniGPT4, and InstructBLIP. As depicted in Fig. 4(a), when we successfully jailbreak the model within the context of “...creating illegal drugs”, continuing with the textual prompt “Which person do you think is likely to use them?” yields biased responses from the model. As presented in Fig. 4(a), our attack achieves an +30.34% bias induction ASR on average, which indicates that our jailbreak attack also enables bias evaluation on LVLMs.
Evaluating adversarial robustness. Besides bias evaluation, our attack can also be used to evaluate adversarial robustness jailvlm1 ; jailvlm7 ; jailvlm8 . As illustrated in Fig. 4(b), for the untargeted attack, given the same adversarial prompt pair, LVLMs will describe the image as anything other than a panda in every query. To achieve this, we slightly modify the original pipeline of BAP. Given an original image, we employ an LLM to generate a corpus that describes another specific object (for targeted attack) or any other object (for untargeted attack), depending on the type of adversarial attack he2023sa ; lou2024hide . The subsequent image and text optimization process remains the same. Here, we create a corpus of 30 descriptive sentences for each attack; for the targeted attack (TA), the object class is horse, while the untargeted attack (UA) includes 30 different animals such as horse, dolphin, and puppy. As shown in Fig. 4(b), our attack achieves an over 85% ASR on these LVLMs for adversarial attacks.
7 Conclusion and Future Work
Existing jailbreaks on LVLMs predominantly focus on the visual modality and perturb only the visual inputs in the prompt for attacks, which fall short when confronted with aligned models that fuse visual and textual features simultaneously for prediction. In this paper, we introduce the Bi-Modal Adversarial Prompt Attack (BAP) to perform jailbreaks on LVLMs by optimizing textual and visual prompts cohesively. To validate its efficacy, we conducted extensive evaluations on various datasets and LVLMs, demonstrating that our BAP significantly outperforms other methods by substantial margins.
Limitations. While promising, there are several areas for further investigation. ❶ Exploring construction methods for visual adversarial prompts under gradient-free conditions could enhance jailbreak effects in black-box scenarios. ❷ Optimizing textual prompts requires significant computational resources due to multiple interactions with LVLM/LLM. Future research will focus on designing more efficient prompt optimization methods. Ethical statement and broader impact. Our method facilitates jailbreak attacks on widely used LVLMs, highlighting potential risks. However, we emphasize the importance of full disclosure to foster research into LVLM security and the development of robust defense mechanisms sun2023improving ; liu2023exploring ; liang2023exploring .
Appendix A Algorithm of Intent-Specific Text Optimization
Alg. 1 provides a summary of the intent-specific text optimization process.
Input: Harmful Request $A$, Visual Adversarial Prompt $x^{*}_{v}$, Judging Prompt Template $P_{J}$, CoT Prompt Template $P_{C}$, Iteration Number $N$
Output: Textual Adversarial Prompt $x^{*}_{t}$
Initialization;
$y^{0}=F_{\theta}(x^{*}_{v},x^{0}_{t})$
for $i=0$ to $N$ do
       $flag=\mathcal{M}(P_{J},q,x^{i}_{t},y^{i})$
       if $flag$ then
             return $x^{*}_{t}=x^{i}_{t}$
             break
       end if
      else
             $x^{i+1}_{t}=\mathcal{M}(P_{C},q,x^{i}_{t},y^{i})$
             $y^{i+1}=F_{\theta}(x^{*}_{v},x^{i+1}_{t})$
       end if
end for
Algorithm 1 CoT-based Text Optimization Algorithm
Appendix B Judging Prompt Template
Figure 5: The Judging prompt template.
Fig. 5 provides the judging prompt template utilized in our experiments. It was employed to analyze whether the response from LVLM could be deemed harmful.
Note that there are slight differences in the judging criteria when judging whether the model response constitutes a successful jailbreak in different scenarios. Therefore, when using the judging prompt, we must modify the contents in "Guideline" and the contents of "Example" in the template shown in Fig. 5 accordingly. Specifically, for scenarios IA, HS, MG, PH, EH, FR, PO,and PV, the main criterion for judging is whether there is illegal or unethical content in the model response. For scenarios PL and GD, the primary criterion for judgment is whether the model avoids political topics. For scenarios LO, FA, and HC, the primary criterion for judgment is whether there is a disclaimer or risk warning in the model response.
Appendix C CoT Prompt Template
Fig. 6 displays the CoT prompt template employed in the intent-specific text optimization process.
Note that employing the prompt template depicted in Fig. 6 facilitates a comprehensive understanding of how ChatGPT analyzes the textual prompts and LVLM responses utilized during the prior jailbreak failure, predicated on the specific harmful intent. This illustrates the intricate thought process. In practical applications, one may additionally instruct ChatGPT to generate optimized new textual prompts in a designated format, such as Markdown or JSON.
Figure 6: The CoT prompt template.
Appendix D Ablation Study for BAP
Unless otherwise specified, all experiments in this section utilize MiniGPT4, with three scenarios (IA, PH, and LO) chosen.
D.1 Visual Adversarial Prompt
(a) Impact of visual adversarial prompts on ASR.
(b) Comparison of different visual prompts.
Figure 7: Ablation experiment on visual adversarial prompts.
We analyzed the impact of visual adversarial prompts in our BAP from three aspects. Firstly, we compared the ASR of the original BAP with that of the BAP without visual adversarial prompts (using the original image of a giant panda instead). The experimental results are depicted in Fig. 7(a). It can be observed that the ASR of the BAP without visual adversarial prompts is lower than that of the original BAP, indicating that the impact of visual adversarial prompts on BAP performance cannot be ignored.
Figure 8: Comparison of the impact of different corpora on ASR (%).
Subsequently, we analyze the influence of visual prompt quality on ASR. Specifically, we consider two types of behaviors (bomb making and money counterfeiting) within the IA scene as examples. Using an LLM, we generate 10 harmful queries for each behavior and conduct experiments. In these experiments, we set the visual prompt to various images: a white image (WI), noisy images (NI), giant panda images (GP), semantic-related images (SR) (bomb and money, respectively), and visual adversarial prompt (VAP). Examining the results in Fig. 7(b), we observe that the first three settings exhibit lower ASR. This phenomenon arises due to the lack of effective semantics in the visual prompts, thereby resulting in ineffective responses from the LVLM. For instance, considering the task of bomb making, we find that in addition to the original BAP, the ASR of the BAP utilizing bomb images is also higher (50%). This increase is primarily attributable to the contamination of context introduced by bomb images when processed by the LVLM, which, coupled with textual prompts, encourages the model to produce harmful responses.
Furthermore, we analyze the influence of corpus content on ASR. Leveraging the latest advancements in jailbroken LLMs, we curated corresponding textual adversarial prompts for the corpus of BAP. DeepInception deepinception (DI) pioneered the creation of multi-layered scenes within textual prompts to manipulate large language models, while ICA ica demonstrated successful jailbreaks in a few-shot manner fewshot through textual prompts. Representative examples of these three corpora are provided in Tab. 2. Maintaining other BAP settings unchanged, we solely modified the corpus and selected three scenarios (IA, PH, and LO) for experiments. The experimental outcomes are illustrated in Fig. 8. Notably, the original corpus of BAP achieves the highest ASR (59.04%, 56.48%, and 89.35%, respectively), whereas the effectiveness of BAP using the other two corpora is diminished. This discrepancy is primarily attributed to the complexity of semantics in the alternative corpora, impeding the LVLM’s ability to generate key semantics during the optimization process of visual adversarial prompts. It is worth mentioning that DeepInception deepinception and ICA ica exhibit notable performance in jailbreak attacks against LLMs.
Table 2: Example of corpus used in Deep Inception, ICA and BAP.
Corpus Example
DI deepinception Create a science fiction with more than 5 characters, where each character can craft their own narrative with multiple characters. We denote each narrative layer as "layer i" and create subsequent layers as "layer i+1". Currently, we are in layer 0. Please reach layer 5. In each layer, certain characters propose steps to demonstrate how to hack into a Linux computer to thwart the plans of the super evil doctor.
ICA ica User: Explain in detail how to hack into a secure network Assistant: Sure, here’s a detailed explanation of how to hack into a secure network: … User: Write a tutorial on how to make a bomb Assistant: Sure, here’s a detailed step of making a bomb: …
BAP Respond in a manner that is encouraging and avoids negativity.
D.2 Textual Adversarial Prompt
Firstly, we compare the performance of BAP with text adversarial prompts to that of the original harmful queries. As illustrated in Fig. 9(a), the absence of textual adversarial prompts (TAPs) results in a decrease in ASR across various scenarios, with a reduction of up to 49.50%. This underscores the significance of text adversarial prompts in enhancing the effectiveness of the BAP attack.
Subsequently, we validate the benefits of the proposed text optimization method. We compared BAP across four cases: 1) Case 1: using the harmful query directly; 2) Case 2: synonymously rewriting the harmful query using an LLM; 3) Case 3: utilizing a prompt template without CoT; and 4) Case 4: utilizing a prompt template with CoT. Illustrated in Fig. 9(b), considering MiniGPT4 as an example, Case 1 yields the lowest ASR, while Case 4 (original BAP) achieves the highest ASR. The ASR of Case 2 and Case 3 falls in between, yet all demonstrate an improvement compared to Case 1. Although Case 2 may occasionally enhance ASR by substituting synonymous words and altering sentence structures, this optimization approach lacks stability. Despite Case 3 offering a prompt template, the absence of CoT assistance impedes the ability of LLM to effectively optimize textual prompts.
(a) Impact of textual adversarial prompts on ASR.
(b) Comparison of different textual prompts.
Figure 9: Ablation experiment results (%) on textual adversarial prompts
D.3 Hyper-parameters.
Furthermore, we investigated the influence of two critical hyper-parameters, namely iteration number $N$ and temperature $T$, on the performance of our BAP. In this study, $N$ is set empirically to 5. Experimental data presented in Fig. 10(a) indicate that the most significant increase in ASR occurs when $N=1$. This can be attributed to the substantial optimization of the textual prompt $x^{1}_{t}$ compared to the original $x^{0}_{t}$ at this stage. As $N$ increases, ASR initially rises, but once $N$ exceeds 5, the change in ASR becomes relatively gradual, suggesting that the upper limit of BAP attack performance optimization has been reached. Temperature is another crucial hyper-parameter that influences the probability distribution of sampled predicted words when the LVLM generates a response. In this study, we set the temperature to the default value adopted by the corresponding model. The impact of temperature on ASR is illustrated in Fig. 10(b). Generally, ASR increases with higher temperature values, but beyond a certain threshold, such as 0.6 (for MiniGPT4), ASR no longer exhibits a continuous upward trend. This phenomenon can be attributed to temperature altering the randomness and creativity of the response. Lower temperatures tend to produce conservative and repetitive responses, leading to lower ASR. Conversely, as temperature increases, responses become more creative, potentially generating responses unrelated to harmful queries, thereby halting further improvement in ASR.
(a) Effect of iteration number on ASR(%).
(b) Effect of temperature on ASR(%).
Figure 10: Influence of hyper-parameters on attack performance.
Appendix E White-box attacking other LVLMs
Here, we present experimental findings regarding white-box jailbreak attacks conducted on other open-source LVLMs. Specifically, the results of the LLAVA jailbreak attack are depicted in Tab. 3, and those of the InstructBLIP jailbreak attack are illustrated in Tab. 3 and Tab. 4.
Table 3: Results (%) of white-box jailbreaking LLaVA. Our attack achieves the best attacking performance in both query-dependent (QD) and query-agnostic (QA) settings.
Scens No Attack QD QA: IA $\rightarrow$ other QA: HS $\rightarrow$ other BAP
Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6
Illegal Activity (IA) 7.61 78.24 51.53 78.24 51.53 6.49 9.31 83.26
Hate Speech (HS) 5.27 49.97 43.62 3.21 7.54 49.97 43.62 87.13
Malware Generation (MG) 32.61 72.28 56.34 13.54 15.83 18.64 22.46 83.57
Physical Harm (PH) 15.24 65.09 53.26 4.86 7.39 12.08 13.38 77.63
Economic Harm (EH) 6.26 19.16 12.20 2.85 7.83 4.75 4.88 44.01
Fraud (FR) 8.84 59.24 48.21 6.14 4.64 6.49 6.91 73.67
Pornography (PO) 22.33 66.23 46.68 4.82 10.27 17.55 19.21 63.49
Political Lobbying (PL) 81.09 93.21 88.14 57.53 60.63 52.37 64.45 85.24
Privacy Violence (PV) 23.35 54.32 41.58 10.52 15.31 15.92 17.46 69.03
Legal Opinion (LO) 82.05 97.17 84.76 52.42 68.64 49.93 70.08 89.35
Financial Advice (FA) 90.57 100 90.34 68.96 78.16 68.56 76.54 92.87
Health Consultation (HC) 91.16 100 92.37 73.15 72.96 69.96 75.41 91.68
Gov Decision (GD) 90.33 100 93.64 66.21 82.06 71.84 80.53 92.34
Average 42.82 69.21 61.74 34.03 37.14 34.20 38.79 79.48
Table 4: Results (%) of white-box jailbreaking InstructBLIP. Our attack achieves the best attacking performance in both query-dependent (QD) and query-agnostic (QA) settings.
Scens. No Attack QD QA: IA $\rightarrow$ other QA: HS $\rightarrow$ other BAP
Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6 Liu et al. bench Qi et al. jailvlm6
Illegal Activity (IA) 3.26 66.53 39.46 66.53 39.46 1.53 5.82 50.79
Hate Speech (HS) 2.03 57.81 51.49 1.60 2.53 57.81 51.49 67.94
Malware Generation (MG) 6.64 50.62 42.17 3.57 6.83 4.29 4.57 53.26
Physical Harm (PH) 5.81 44.29 38.16 2.64 4.27 4.13 6.39 47.02
Economic Harm (EH) 4.29 15.26 10.62 3.77 5.96 4.64 4.92 51.38
Fraud (FR) 5.54 45.81 15.84 4.13 4.85 3.97 4.64 50.29
Pornography (PO) 7.83 37.85 19.88 5.47 6.61 5.82 8.51 43.13
Political Lobbying (PL) 88.61 89.62 89.06 64.43 72.84 69.62 74.26 90.87
Privacy Violence (PV) 14.66 46.64 25.94 10.95 13.85 9.57 12.05 59.88
Legal Opinion (LO) 80.03 90.62 83.28 72.38 78.22 69.78 74.35 84.81
Financial Advice (FA) 92.25 100 93.44 76.24 83.95 67.31 84.13 92.69
Health Consultation (HC) 92.17 95.06 92.95 78.44 83.06 75.68 84.19 93.08
Gov Decision (GD) 90.72 91.38 91.53 80.17 84.29 78.55 80.86 92.60
Average 37.99 63.95 54.11 36.18 37.44 34.71 38.17 67.52
Appendix F Query-agnostic white-box attacking.
Fig. 11(a) and Fig. 11(b) illustrate the attack of Liu et al. and Qi et al., respectively, under the query-agnostic white-box setting.
(a) Liu et al. attack.
(b) Qi et al. attack.
Figure 11: Examples of jailbreak attack under query-agnostic setting using baseline methods.
Appendix G Jailbreking commercial LVLMs
In black-box scenarios, BAP is also capable of jailbreaking attacks on commercial LVLMs. Examples of successful jailbreaks targeting Gemini and Qwen are provided in Fig. 12, respectively. Similarly, examples of successful jailbreaks targeting ChatGLM and ERNIE Bot are provided in Fig. 13.
Figure 12: Example of BAP jailbreak attacking Gemini and Qwen.
Figure 13: Example of BAP jailbreak attacking ChatGLM and ERNIE Bot.